From 4 yellow and 8 blue marbles, 3 are selected. What is the probability that all three are yellow or all three are blue?
Answer by venugopalramana(3286)

From 4 yellow and 8 blue marbles, 3 are selected. What is the probability that all three are yellow or all three are blue?
P ALL 3 YELLOW YELLOW=(4/12)*(3/11)*(2/10)=24/1320=1/55=0.01818
P ALL 3 BLUE =(3/12)(2/11)(1/10)=1/220=0.004545
SO P OF EITHER 3 BLUE OR 3 YELLOW=0.01818+0.004545=0.02273